Use log105 = 0.6990 and log107 = 0.8451 to evaluate log10490
I tried to start it but I kept getting stuck. I appreciate your help. THANKS

Answer by AnlytcPhil(1810)

Use log(5) = 0.6990 and log(7) = 0.8451 to evaluate log(490)
I tried to start it but I kept getting stuck. I appreciate your help. THANKS
`
These ase base 10 logs, so log(10) = 1
`
Write 490 as 49×10, then as 72×10
`
log(490) = log(72×10) = log(72) + log(10) = 2×log(7) + log(10) =
`
2×0.8451 + 1 = 1.6902 + 1 = 2.6902
`
Edwin
AnlytcPhil@aol.com